OpenStudy (anonymous):

How to solve: Use geometric series 1/(x+1) = Summation (n=0 to infinity) (-1)^n * x^n, to find the power series representation of f(x) = 1 / (4x +3).

OpenStudy (anonymous):

F(x)=1/(x+1) can be represented as \[\sum_{0}^{infiniti}(-1)^n x^n\] F(x)=1/(4/3x +1) is same as f(x)=1/4x+3 So you would substitute in 4/3x in place of x

OpenStudy (anonymous):

Does it at least make some sense?

OpenStudy (anonymous):

You would have to multiply the summation by 3.

OpenStudy (anonymous):

1/3(4/3x+1) you can factor out 1/3

OpenStudy (anonymous):

* actually 1/3

OpenStudy (anonymous):

Yes you factor out the 1/3 but you did not account for it.

OpenStudy (anonymous):

I missed that

OpenStudy (anonymous):

it is one of those little errors.

OpenStudy (anonymous):

So Does it make sense now WRosa?

OpenStudy (anonymous):

\[1/3\sum_{0}^{\infty} (-1^n)(4/3x)^n\]

OpenStudy (anonymous):

So the solution is: f(x) = 1/(4x+3) f(x) = 1/3 ((4/3)x + 1) \[1/3 \sum_{n=0}^{\infty} (-1)^{n} ((4/3)x)^{n}\]